A 500 gram, 8 centimeter diameter can is filled with uniform dense food. It rolls across the floor at 1 m/s. Find the kinetic energy.
Answer #1

Answer question
Answer #2

Can's Kinetic Energy = Translational KE + Rotational KE
KE of can = (1/2)mv^2 + (1/2)I?^2
KE of can = (1/2)mv^2 + (1/2)[(1/2)(m)r^2]?^2
KEof can=(1/2)(0.500)(1^2)+(1/2)[(1/2)(0.500)...

v = ?r
? = v/r = 1/0.04 = 25 rad/sec

KEcan = (1/2)(0.500)(1^2)+(1/2)[(1/2)(0.500)(0.0...
= 0.312 J ANSWER
Answer #3

Can's Kinetic Energy = Translational KE + Rotational KE
KE of can = (1/2)mv^2 + (1/2) * I * ?^2
Solid cylinder has I=.5*m*r^2

r = 8/2 = 4 cm = 0.04 m
v = ?r
? = v/r = 1/0.04 = 25 rad/sec

I = 0.5 * .500*0.04*0.04 = 0.0004


KEcan = (1/2) * (0.500)* (1*1) + 0.5* 0.0004 * 25 * 25
= 0.25 + 0.125 = 0.375